OpenStudy (anonymous):

If cosθ= -15/13 and sin θ< 0, what is tan θ?? a.12/5 b. 13/5 c. -12/5 d. sqrt 194/5

OpenStudy (kc_kennylau):

Use \(\cos^2\theta+\sin^2\theta=1\) to find \(\sin\theta\) first

OpenStudy (anonymous):

in the first theta i put -15/13? and the second i put (theta)<0?

OpenStudy (kc_kennylau):

\(\cos\theta\) is NOT \(\cos\) times \(\theta\)

OpenStudy (kc_kennylau):

\(\cos\theta\) is one entity; \(\sin\theta\) is one entity

OpenStudy (anonymous):

i dont understand im sorry im trying

OpenStudy (kc_kennylau):

Treat \(\cos\theta\) as one thing instead of two things

OpenStudy (kc_kennylau):

And bear in mind the notation: \(\cos^2\theta=(\cos\theta)^2\)

OpenStudy (kc_kennylau):

So the formula would be \((\cos\theta)^2+(\sin\theta)^2=1\)

OpenStudy (kc_kennylau):

Substitute \(\cos\theta=-\dfrac{15}{13}\) inside to find the value of \(\sin\theta\)

OpenStudy (anonymous):

i got no solution -.-

OpenStudy (kc_kennylau):

Oh, sorry, I forgot to check the validity

OpenStudy (kc_kennylau):

\(\cos\theta\) and \(\sin\theta\) must be between -1 and 1

OpenStudy (kc_kennylau):

So no solution

OpenStudy (anonymous):

so noww what do i do to find tan(theta)

OpenStudy (kc_kennylau):

You can't; theta doesn't even exist

OpenStudy (anonymous):

so then what would the answer be? :0

OpenStudy (anonymous):

13/5?

OpenStudy (kc_kennylau):

An answer doesn't exist

OpenStudy (anonymous):

i found out it was 12/5 idk how

OpenStudy (kc_kennylau):

No it isn't; it doesn't even exist

OpenStudy (kc_kennylau):

Or did you mistype the question?

OpenStudy (anonymous):

nope my school just dont make sence